# A Semantic Search Engine for Mathlib4

### v1

@@ 4.3 Performance Measures
To compare different retrieval paradigms’ performance on our labeled dataset, three commonly used metrics are adopted in our benchmark. Precision@k calculates the average relevant document proportion in top k retrieved results among the query set :
where stands for the retrieved theorem list for i-th query, and if and only if the j-th retrieved result of i-th query is "Exact Match", otherwise . Based on the assumption that all unlabeled theorems are irrelevant, it is also reasonable to calculate the Recall@k:
@@
where stands for the score of j-th retrieved theorem with respect to i-th query, defined in Table 2. We note that our score is not proportional to label number, aiming to emphasize the importance of exact matching in mathlib4 retrieval task. We further define as the highest possible by properly arranging retrieved theorems, and finally:
To sum up, nDCG@k measure both the retrieving and ranking ability of the given engine in a unified and detailed way. Meanwhile, P@10 (Precision@10) and R@10 (Recall@10) focus on the retrieval effectiveness solely. In our benchmark, all three metrics will be reported both on the whole query set and on each query category.
@@ 5 Experiments
In this section, we evaluate the performance of various theorem retrieval methods using our benchmark. The experimental configurations, including embedding models, task instructions and computational expenses, are detailed in Section 5.1. We compare the performances of different retrieval methodologies in Section 5.2. Section 5.3 analyzes the impact of different task instructions, the documents content type, and the query augmentations employed in our approach.
@@ 5.1 Experiment setup
We have considered four embedding models in our experiments: text-embedding-ada-002 and text-embedding-3-large from OpenAI, UAE-Large-V1[13], and [31]. We use default task instructions in UAE-Large-V1. For , asymmetrical task instructions were employed in the non-augmented query setting, and symmetrical task instructions were applied in the query augmentation setting, as shown in Table 3. We observe that the original implementation of does not use document-side task instructions to reduce computational cost during document indexing in multiple retrieval tasks. Considering the complexity of the mathematical embedding task, we have modified this to a two-sided prompt approach, as has shown substantial performance degradation with one-sided prompt setting in our benchmarks.
| Query aug. & Doc Type | Side | Input with Task Instructions |
None & Formal |
Query |
[…]
Doc |
[…]
F+IF & F+IF |
Query |
[…]
Doc |
[…]
For the baseline models, Moogle, along with all four embedding models applied on original Lean 4 formal corpus and unaugmented query, are used in our experiments. The same four embedding models equipped with formal + informal query augmentation and formal + informal document corpus are also tested on our benchmark to demonstrate the efficacy of our approach. During evaluation, inputs to the embedding model exceeding 4096 characters were truncated due to GPU memory limitations. For all tested models, the corpus embedding process was completed within three hours on a single V100 GPU. Subsequently, evaluations of these models were conducted using the three metrics mentioned earlier.
@@ 5.2 Main Results
Table 4 and 5 present the results of Moogle and four embedding models. , when integrated with our retrieval pipeline, achieves the best performance across three overall metrics and significantly outperforms all other methods, including its own performance on a formal corpus with unaugmented queries. We observe that all retrieval methods benefit from using an augmented corpus and queries, as augmentation expands concrete mathematics and Lean 4 terms, and these embedding models typically perform better on symmetrical retrieval tasks [31].
Examining the results averaged by category, we find that attains the best results in nearly all categories, with the exception of nDCG@20 for Lean 4 Terms. In this case, the augmented informal information may impair retrieval performance, as it is not essential in formal-formal retrieval, which focuses on lexical matches, but it does alter the final embedding representation. Meanwhile, our method markedly enhances performance in the Theorem Name category, because the expansion of the statement for a given theorem name is crucial for successful retrieval in this category. Our approach of informalization and query augmentation facilitates this process for embedding models.
@@ 5.3 Ablation Studies
We demonstrate that our pipeline enables embedding models to unpack concrete mathematical concepts, transforming the retrieval task into a symmetric one and thereby enhancing their performance in Mathlib4 retrieval tasks. In this subsection, we analyze our results and conduct ablation studies from various perspectives. The necessity of proper task instructions is illustrated in Section 5.3.1, while the impact of document preparations and query augmentation is examined in Sections 5.3.2 and 5.3.3, respectively.
5.3.1 Necessity of proper task instructions
Recent models incorporate task instructions into their training to enhance zero-shot performance [28, 31]. In this part, we analyze the impact of task instructions in mathematical informal retrieval. We present a visualization of the embeddings of all 50 queries, generated by four embedding engine in our benchmark, using t-SNE[29]. We use default task instruction for UAE-Large-V1. For , in addition to using empty instruction, we also test mathematics retrieving instruction "Given a math search query, retrieve theorems mathematically equivalent to the query" and Lean 4 retrieving instruction "Given a math search query, retrieve Lean 4 written theorems mathematically match the query".
As illustrated in Figure 5, the cluster of theorem names appear in the results produced by all four embedding models. However, these clusters vanish when we provide with mathematics-aware task instructions. Given that we assume all queries within the same query group (denoted by identical colors in Figure 5) as having the same search intent, they should be proximal on the graph, as t-SNE maintains the relative distance relationships between vectors. Thus, the more effective the embedding engine, the more closely the dots of the same color group together. A comparison of subfigures 4(c) and 4(d) with 4(a) and 4(b) reveals that text-embedding-3-large and demonstrate superior performance in discerning mathematical search intentions, aligning with our main results (Table 4).
The presence of theorem name clusters in these subfigures, however, contradicts this principle, suggesting that the embedding models perceive these theorem names as more similar to each other rather than correctly associating them with their respective query groups. While this might be acceptable for other retrieval tasks, it is clearly inappropriate for our mathematical information retrieval context. In contrast, the absence of theorem clusters in subfigures 4(e) and 4(f) suggests that these embedding methods successfully group theorem names with their corresponding statements, despite their notable differences in appearance and structure. This indicates their proficiency in comprehending mathematical theorems and recognizing search intents. In summary, appropriate task instructions enhance the embedding models’ sensitivity to the search intents of mathematical concepts, thereby improving retrieval effectiveness on Mathlib4.
@@ 5.3.2 Ablation of document preparations
None & Informal |
Query |
Given a math search query, retrieve theorems mathematically equivalent to the query |
Doc |
Represent the given math theorem statement for retrieving related statement by natural language query | |
None & F+IF |
Query |
Given a math search query, retrieve theorems stated in bilingual Lean 4 + natural language that mathematically match the query |
Doc |
Represent the given formal math statement written in Lean 4 concatenated with its natural language explanation for retrieving related statement by natural language query | |
Formal & Formal |
Query & Doc |
Retrieve math theorems stated in Lean 4 that are mathematically equivalent to the given one |
IF & IF |
Query & Doc |
Retrieve math theorems that are mathematically equivalent to the given one |
Table 6 presents the effects of changing corpus type with or without query augmentation. We simply use the original Lean 4 statement and the informalized statement in our formal + informal corpus to form the formal and informal corpus. Under no query augmentation setting, results indicate that both engines underperform with incomplete corpus components, yielding lower overall scores on all metrics. Specifically, the formal + informal corpus enhances performance on Lean 4 Terms by incorporating formal data and improves score on the Natural Description category by providing hybrid information. For the Theorem Name category, significantly benefits from the formal + informal corpus by utilizing the unfolded mathematical descriptions provided by the informalized statement, whereas text-embedding-3-large shows diminished results, likely due to the lack of training on mixed-domain texts and absence of adaptive instructions. We note that formal corpus outperforms formal+informal corpus on Lean 4 Terms category for , with the same reason we explained in Section 5.2: retrieval on Lean 4 Terms highly relies on lexical instead of semantic information, and the existence of informal information alters the final embedding vector. Similar results can be observed when using augmented queries.
When using non-augmented queries, a more detailed examination of the results that takes averages on query groups, further validates our analysis. As depicted in Figure 6, in the majority of instances, the "formal + informal" configuration does not experience the abrupt decline in performance observed in solely formal or informal settings. The informalized statement provided in the corpus mitigate the lexical mismatch between query and formal statement. Furthermore, the incorporation of both informal and formal information occasionally leads to enhanced outcomes. These two arguments support the choice of hybrid corpus in our pipeline.
@@ 5.3.3 Ablation of query augmentations
The effect of query augmentation can be observed by comparing the results in the upper and lower sections of Table 6. Query augmentation is designed to align the types of queries and documents; thus, we use the formal and informal part of our augmented queries for the formal and informal corpora, respectively. As indicated in the table, augmented queries enhance the performance of all methods across the board. Specifically, query augmentation markedly improves the performance of all methods in identifying Theorem Names by expanding terms to provide a richer context. We claim that the process of informalization serves a similar purpose, and that such additional information can be extracted through the use of appropriate instructions. This is evidenced by the high nDCG@20 scores achieved by on Theorem Names by using informal or formal + informal corpus, under the setting of using non-augmented query. The comparable overall performance of on formal + informal corpus in this case, presents it as a cost-effective option for frequent usage due to its potential to reduce LLM API costs.

### v2

@@ 4.3 Performance Measures
To compare different retrieval paradigms’ performance on our labeled dataset, three commonly used metrics are adopted in our benchmark. Precision@k calculates the average relevant document proportion in top- retrieved results among the query set :
where represents the list of retrieved theorems for the -th query, and if and only if the -th retrieved result of -th query is "Exact Match", otherwise . Assuming that all unlabeled theorems are irrelevant, we can also calculate Recall@k:
@@
where represents the score of the -th retrieved theorem for the -th query, as defined in Table 2. Our scoring system is not proportional to the label number, emphasizing the importance of exact matching in the mathlib4 retrieval task. We further define as the highest possible by properly arranging the retrieved theorems, and finally:
To sum up, nDCG@k measures both the retrieval and ranking ability of the given engine in a unified and detailed way. Meanwhile, P@10 (Precision@10) and R@10 (Recall@10) focus solely on the retrieval effectiveness. In our benchmark, all three metrics will be reported for both the entire query set and each query category.
@@ 5 Experiments
In this section, we evaluate the performance of various theorem retrieval methods using our benchmark. The experimental setup is detailed in Section 5.1. We compare the performances of different retrieval methods in Section 5.2. Additionally, the ablation study on document preparations and query augmentations is presented in Section 5.3.
@@ 5.1 Experiment Setup
We have considered four embedding models in our experiments: text-embedding-ada-002 and text-embedding-3-large from OpenAI, UAE-Large-V1(Li and Li, 2023), and (Wang et al., 2024). For the baseline methods, we compare BM25 (Robertson and Zaragoza, 2009), Moogle, and the four embedding models applied to the original Lean 4 formal corpus and unaugmented query. The same four embedding models equipped with formal + informal query augmentation and formal + informal document corpus are also tested on our benchmark to demonstrate the efficacy of our approach. The implementation details are provided in Appendix C.
@@ 5.2 Main Results
Table 3 and 4 present the results of BM25, Moogle, and four embedding models. The model , when integrated with our retrieval pipeline, achieves the best performance across three overall metrics and significantly outperforms all other methods, including its own performance on a formal corpus with unaugmented queries. We observe that all retrieval methods benefit from using an augmented corpus and queries, as augmentation expands concrete mathematics and Lean 4 terms, and these embedding models typically perform better on symmetrical retrieval tasks (Wang et al., 2024).
Examining the results averaged by category, we find that attains the best results in nearly all categories, with the exception of nDCG@20 for Lean 4 Terms. This exception is likely due to the augmented informal information negatively impacting retrieval performance, as it is not essential in formal-formal retrieval (formal queries, formal results) which focuses on lexical matches. Notably, our method significantly improves performance in the Theorem Name category. The enhancement is due to the expansion of the statement for a given theorem name by our query augmentation, which is essential for successful retrieval in this category. Furthermore, the combined corpus of formal and informal statements facilitates the retrieval process.
@@ 5.3 Ablation Studies
In this subsection, we perform ablation studies on document preparation and query augmentation. The results of these studies are presented in Sections 5.3.1 and 5.3.2, respectively. Additionally, the ablation study on task instructions in provided in Appendix D.
@@ 5.3.1 Ablation of Document Preparations
Table 5 presents the effects of changing corpus type with and without query augmentation. We use the original Lean 4 statement and its informalized version to create the formal and informal corpus. Without query augmentation, results indicate that both OpenAI v3 large and underperform with incomplete corpus components, yielding lower overall scores on all metrics. Specifically, the formal + informal corpus enhances performance on Lean 4 Terms by incorporating formal data and improves score on the Natural Description category by providing hybrid information. For the Theorem Name category, significantly benefits from the formal + informal corpus by utilizing the unfolded mathematical descriptions provided by the informalized statement, whereas OpenAI v3 large shows diminished results, likely due to the lack of training on mixed-domain texts and absence of adaptive instructions. We note that the formal corpus outperforms the formal+informal corpus in the Lean 4 Terms category for , as explained in Section 5.2: retrieval on Lean 4 Terms relies heavily on lexical rather than semantic information, and the presence of informal information negatively impacts retrieval performance. Similar results are observed with augmented queries.
A more detailed examination of the results for each query group without query augmentation is shown in Figure 4. In most instances, the "formal + informal" configuration avoids the abrupt performance decline observed in solely formal or informal settings. The informalized statement provided in the corpus mitigates the lexical mismatch between the query and the formal statement, which justifies our choice of a combined corpus.
@@
| None & Informal | Query | Given a math search query, retrieve theorems mathematically equivalent to the query |
| Doc | Represent the given math theorem statement for retrieving related statement by natural language query | |
| None & F+IF | Query | Given a math search query, retrieve theorems stated in bilingual Lean 4 + natural language that mathematically match the query |
| Doc | Represent the given formal math statement written in Lean 4 concatenated with its natural language explanation for retrieving related statement by natural language query | |
| Formal & Formal | Query & Doc | Retrieve math theorems stated in Lean 4 that are mathematically equivalent to the given one |
| IF & IF | Query & Doc | Retrieve math theorems that are mathematically equivalent to the given one |
@@ 5.3.2 Ablation of Query Augmentations
The effect of query augmentation is evident when comparing the results in the upper and lower sections of Table 5. As shown, augmented queries enhance the performance across all compared methods. Specifically, query augmentation significantly improves the ability of all methods to identify Theorem Names by expanding terms and providing richer context. Notably, the performance of on the formal + informal corpus without query augmentation is comparable to its performance on the same corpus with query augmentation. This makes it a cost-effective option for deploying our semantic search engine, as it eliminates the need to prompt an LLM with every query.
@@ 6 Conclusion
7 Limitations
When translating mathlib4 theorems into their informal versions, we extract the related definitions through the hyperlinks in the theorem statements. This approach to extracting dependencies can sometimes be inaccurate and may omit certain dependencies. A more precise method is to extract the dependencies of a theorem by interacting with Lean’s language server.
Another limitation of our work is the use of , a 7B embedding model, in our search engine. This model may pose challenges when deploying the search engine on resource-constrained devices like laptops. In future work, we aim to improve the performance of smaller embedding models for theorem retrieval tasks to enable broader accessibility.
Acknowledgment
This work is supported in part by the New Cornerstone Investigator Program.
Appendix A Example of Informalized Mathlib4 Statements
In this section, we provide an example comparing the informalized mathlib4 statement generated by (Jiang et al., 2023), which used GPT-4 without any additional context, to the statement produced by our approach that incorporates additional context.
Formal Statement:
theorem dite_eq_iff { : Sort u_2} {P : Prop} {c : } {A : P → } {B : P → } : dite P A B = c ( h, A h = c) h, B h = c
Informal Statement (Jiang et al., 2023)111111https://github.com/albertqjiang/MMA/tree/main/data:
dite P A B equals c if and only if there exists a proof h such that A h equals c, or there exists a proof h such that B h equals c.
Informal Statement (ours):
Dependent if-then-else equivalence: For any proposition P that is decidable, and any elements c, A, and B, the expression ‘if P then A else B’ is equal to c if and only if either there exists a proof h such that A h is equal to c, or there exists a proof h such that B h is equal to c.
In this example, the informalized statement provided by (Jiang et al., 2023) retains the "dite" notation, which may be difficult for readers unfamiliar with mathlib4 naming conventions. In contrast, our informalized statement translates "dite" into "dependent if-then-else," making the statement more accessible. This translation was achieved by extracting the definition of "dite" and its documentation string "Dependent if-then-else…" in the prompt, demonstrating how additional context enhances the interpretation and conversion of formal theorems into informal language.
Appendix B Prompt for Query Augmentation
In this section, we present the prompt template for query augmentation, as illustrated in Figure 5. We outline specific principles for query augmentation, highlighting the importance of providing precise information, using LaTeX for mathematical expressions, and clarifying ambiguous inputs. Additionally, we offer examples of query augmentations to enhance understanding of the task.
Appendix C Implementation Details
In this section, we provide the implementation details of our experiments.
We use default task instructions in UAE-Large-V1. For , we employ different task instructions for the document and query in the non-augmented query setting and the same task instructions for both in the query augmentation setting, as shown in Table 7. Notably, the original implementation of omits document-side task instructions to reduce computational costs during document indexing in multiple retrieval tasks. Given the complexity of the theorem retrieval task, we have modified this approach to use two-sided prompts.
During evaluation, inputs to the embedding model exceeding 4096 characters were truncated due to GPU memory limitations. For all tested models, the corpus embedding process was completed within three hours on a single Nvidia V100 GPU with 16GB of GPU memory.
| Query aug. & Doc Type | Side | Input with Task Instructions |
[…]
Appendix D Ablation Study on Task Instructions
In this section, we investigate the impact of task instructions on the theorem retrieval performance. We present a visualization of the embeddings of all 50 queries, generated by four embedding engines in our benchmark, using t-SNE(van der Maaten and Hinton, 2008). We use the default task instruction for UAE-Large-V1. For , we test three instructions: an empty instruction, a mathematics retrieval instruction ("Given a math search query, retrieve theorems mathematically equivalent to the query"), and a Lean 4 retrieval instruction ("Given a math search query, retrieve Lean 4 written theorems that mathematically match the query").
As illustrated in Figure 6, clusters of theorem names appear in the results produced by all four embedding models. However, these clusters vanish when we provide with mathematics-aware task instructions. Given that we assume all queries within the same query group (denoted by identical colors in Figure 6) have the same search intent, they should be proximal on the graph, as t-SNE maintains the relative distance relationships between vectors. Thus, the more effective the embedding engine, the more closely the dots of the same color group together. A comparison of subfigures 6(c) and 6(d) with 6(a) and 6(b) reveals that text-embedding-3-large and demonstrate superior performance in discerning mathematical search intentions, aligning with our main results (Table 3).
The presence of theorem name clusters in these subfigures, however, contradicts this principle, suggesting that the embedding models perceive these theorem names as more similar to each other rather than correctly associating them with their respective query groups. While this might be acceptable for other retrieval tasks, it is inappropriate for our mathematical information retrieval context. In contrast, the absence of theorem clusters in subfigures 6(e) and 6(f) suggests that these embedding methods successfully group theorem names with their corresponding statements, despite their notable differences in appearance and structure. This indicates their proficiency in comprehending mathematical theorems and recognizing search intents. In summary, appropriate task instructions enhance the embedding models’ sensitivity to the search intents of mathematical concepts, thereby improving retrieval effectiveness on mathlib4.
Appendix E Models for Query Augmentation
In this section, we compare the performance of two LLMs, DeepSeek-V2-Chat and GPT-4, for query augmentation. Table 8 presents the performance of both models using the informal+formal corpus setting, with results averaged over 10 runs.
| API | nDCG@20 | P@10 | R@10 | Cost per 1M input tokens | Cost per 1M output tokens |
| gpt-4-0125-preview | 0.721 | 0.195 | 0.911 | $10.00 | $30.00 |
| deepseek-chat | 0.749 | 0.202 | 0.933 | $0.14 | $0.14 |
As shown, DeepSeek-V2-Chat not only outperforms GPT-4 in terms of effectiveness but also offers a significantly lower cost. The results suggest that DeepSeek-V2-Chat is a more efficient and cost-effective choice for query augmentation tasks.